# Singularity-free cosmology from interactions in the dark sector

### v1

@@ A.2 Improper Node
Figure 21: The $Z$-$X$ phase space referred to in Appendix A.2. Here, the parameters are set to $w_{x}=-0.6$ and $q=0.847386$. In this case, the stability of $dS_{1+}$ (orange) is an improper node, and the zero acceleration curve intersects the repellor-saddle separatrix twice. Trajectories to the left of the separatrices expand to a singularity, some with and some without a decelerated period. Those to the right of the separatrices all have a decelerated period and accelerate as they asymptotically expand towards $dS_{2+}$ (magenta). Therefore, the trajectories that evolve to the right of the separatrices are qualitatively of interest.
@@
Fig.s 22 and 23 show the 3-D $X$-$Y$-$Z$ phase spaces corresponding to a trajectory that evolves to the right of the separatrices in Fig. 21. Qualitatively, these phase spaces are similar to that of Fig.s 19 and 20. Trajectories with positive spatial curvature are shown in Fig. 22. Two Einstein points exist in the phase space, one of which is a center, and the other is a saddle which forms part of the CFS. The trajectories that evolve outside the CFS are qualitatively interesting, as they contract from $dS_{2-}$, go through a quasi-de Sitter bounce, and then expand towards $dS_{2+}$, intersecting the zero acceleration surface twice during expansion. Expanding flat and open trajectories are shown in Fig. 23 and are all of interest, as they emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and intersect the zero acceleration surface twice. This means they have a decelerated phase, followed by a late-time acceleration as they expand towards $dS_{2+}$.
@@
Figure 22: The $X$-$Y$-$Z$ phase space corresponding to Fig. 21, where $w_{x}=-0.6$ and $q=0.847386$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 21 that evolves to the right of the separatrices. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$ and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between between $dS_{2-}$ and $dS_{2+}$, and intersect the zero acceleration surface twice during expansion, meaning they have a decelerated phase and a late-time acceleration.
@@
Figure 23: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 21, where $w_{x}=-0.6$ and $q=0.847386$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 21 that evolves to the right of the separatrices, expanding to $dS_{2+}$. Both the flat and open trajectories intersect the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and finally have an accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@ Appendix B $q>3$
The $z$-coordinate of $dS_{1+}$ is proportional to $1/q^{2}$ and the $x$-coordinate is proportional to $1/q$. When $q<3$, $dS_{1+}$ will therefore have larger $z$- and $x$-values than when $q>3$. In particular, when $q>3$, $dS_{1+}$ has an $x$-value of $x<1$. This means the repellor-saddle separatrix will be shorter when $q>3$, as the repelling fixed point is closer to the saddle $dS_{3+}$. We find that when $dS_{1+}$ is a spiral repellor, the separatrix and zero acceleration curve can intersect as the vector field spirals out from $dS_{1+}$, which means the separatrix also spirals out from $dS_{1+}$. However, when $dS_{1+}$ is a repelling or an improper node, the separatrix does not spiral out, and so does not intersect the zero acceleration curve. In the following, we present the 2-D and 3-D phase spaces where $dS_{1+}$ is a repelling node and an improper node. The color scheme of the fixed points and critical points in the 2-D $Z$-$X$ phase spaces is as in Table 4, and the color scheme of the curves is given in Table 5. The color scheme of the fixed points and critical points in the 3-D phase spaces is given in Table 8, and the color scheme of the curves and surfaces is as in Table 9.
@@ B.1 Repelling Node
Fig. 24 shows the $Z$-$X$ phase space when $q>3$ and $dS_{1+}$ is a repelling node. All trajectories are non-singular, and expand from $dS_{1+}$ at high energy, and tend towards $dS_{2+}$ at late times. The trajectories that do not intersect the red zero acceleration curve always accelerate, and those that cross the red curve twice have a period of deceleration, followed by a late-time acceleration. There is also a limiting case where a trajectory touches the red curve and has a point of zero acceleration, but otherwise accelerates. The trajectories that intersect the zero acceleration curve twice are the models of interest.
@@
Figure 24: The $Z$-$X$ phase space referred to in Appendix B.1. Here, the parameters are set to $w_{x}=-0.5$ and $q=5.2$. The stability of $dS_{1+}$ (orange) is a repelling node, and the zero acceleration curve does not intersect the repellor-saddle separatrix. Apart from the separatrix itself which evolves to $dS_{3+}$ (dark blue), all other trajectories expand towards $dS_{2+}$ (magenta) at late-times. Trajectories which do not intersect the $a^{\prime\prime}=0$ curve always accelerate, and those that intersect it twice have a decelerated period. One trajectory in the phase space will touch the zero-acceleration curve, but will otherwise accelerate.
@@
Fig.s 25 and 26 show a first integral surface in the $X$-$Y$-$Z$ phase space corresponding to a trajectory that intersects the zero acceleration curve twice in Fig. 24. Fig. 25 shows the trajectories with positive spatial curvature, and Fig. 26 shows expanding flat and open trajectories. All the trajectories in Fig. 25 bounce at least once. Two Einstein points exist at $Y=0$, one of which is a saddle which the CFS passes through, and the other is a center. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, or are cyclic around the center Einstein fixed point. These trajectories are not of interest, as the bouncing trajectories always accelerate, and the cyclic trajectories have a decelerated phase during expansion, but no late-time acceleration. The trajectories that evolve outside the CFS bounce once during a quasi-de Sitter phase, contracting from $dS_{2-}$ then expanding towards $dS_{2+}$. These trajectories intersect the zero acceleration surface twice during expansion, meaning they accelerate though the bounce, becoming flatter as they start to expand, then have a decelerated phase where large scale structure could form, and then have a final late-time accelerated expansion as they approach a late-time cosmological constant represented by $dS_{2+}$. The trajectories outside the CFS are therefore of interest.
@@
Figure 25: The $X$-$Y$-$Z$ phase space corresponding to Fig. 24, where $w_{x}=-0.5$ and $q=5.2$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.4$, which corresponds to a trajectory in Fig. 24 that intersects the red zero acceleration curve twice. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and intersect the zero acceleration surface twice during expansion, meaning they have a decelerated phase followed by a late-time acceleration.
@@
Figure 26: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 24, where $w_{x}=-0.5$ and $q=5.2$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.4$, which corresponds to a trajectory in Fig. 24 that intersects the red zero acceleration surface twice. Both the flat and open trajectories cross the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@ B.2 Improper Node
Figure 27: The $Z$-$X$ phase space referred to in Appendix B.2. Here, the parameters are set to $w_{x}=-0.3$ and $q=3.32488$. Here, the stability of $dS_{1+}$ (orange) is an improper node, and the zero acceleration curve does not intersect the repellor-saddle separatrix. Except for the separatrix which expands to $dS_{3+}$ (dark blue), all other trajectories expand towards $dS_{2+}$ (magenta) at late-times. Trajectories which do not intersect the $a^{\prime\prime}=0$ curve always accelerate, and those that intersect it twice have a decelerated phase. One trajectory in the phase space will touch the zero-acceleration curve, but will otherwise accelerate.
@@
Fig.s 28 and 29 show the 3-D $X$-$Y$-$Z$ phase spaces corresponding to a trajectory that intersects the zero acceleration curve twice in Fig. 27. Qualitatively, the phase spaces are similar to that of Fig.s 25 and 26. Fig. 28 shows the trajectories with positive spatial curvature. Two Einstein fixed points exist in the phase space, one of which is a center and the other is a saddle which forms part of the CFS. The trajectories that evolve outside the CFS and bounce once are the models that are qualitatively interesting. These trajectories contract from $dS_{2-}$, then bounce during a quasi de-Sitter phase, and finally expand towards $dS_{2+}$. They cross the zero acceleration surface twice during expansion, meaning they accelerate through the bounce, becoming flatter as they start to expand, then have a decelerated period followed by a late-time accelerated expansion. Expanding flat and open trajectories are shown in Fig. 29, and are all of interest. Flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and intersect the $a^{\prime\prime}=0$ surface twice. This means they have a decelerated phase where large scale structure could form, and accelerate at late times as they expand towards $dS_{2+}$.
@@
Figure 28: The $X$-$Y$-$Z$ phase space corresponding to Fig. 27, where $w_{x}=-0.3$ and $q=3.32488$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.05$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 27 that intersects the red zero acceleration curve twice. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and intersect the zero acceleration surface twice during expansion, meaning they have a decelerated phase followed by a late-time acceleration.
@@
Figure 29: The expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 27, where $w_{x}=-0.3$ and $q=3.32488$. The trajectories are plotted on a first integral surface where $X_{0}=0.05$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 27 that intersects the red zero acceleration curve twice. Both the flat and open trajectories intersect the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.

### v2

@@ A.2 Improper node
Figure 21: $Z$-$X$ phase space referred to in Appendix A.2. Here, the parameters are set to $w_{x}=-0.6$ and $q=0.847386$. In this case, the stability of $dS_{1+}$ (orange) is an improper node, and the zero acceleration curve intersects the repellor-saddle separatrix twice. Trajectories to the left of the separatrices expand to a singularity, some with and some without a decelerated period. Those to the right of the separatrices all have a decelerated period and accelerate as they asymptotically expand towards $dS_{2+}$ (magenta). Therefore, the trajectories that evolve to the right of the separatrices are qualitatively of interest.
@@
Figures 22 and 23 show the three-dimensional $X$-$Y$-$Z$ phase spaces corresponding to a trajectory that evolves to the right of the separatrices in Fig. 21. Qualitatively, these phase spaces are similar to that of Figs. 19 and 20. Trajectories with positive spatial curvature are shown in Fig. 22. Two Einstein points exist in the phase space, one of which is a center, and the other is a saddle which forms part of the CFS. The trajectories that evolve outside the CFS are qualitatively interesting, as they contract from $dS_{2-}$, go through a quasi-de Sitter bounce, and then expand towards $dS_{2+}$, intersecting the zero acceleration surface twice during expansion. Expanding flat and open trajectories are shown in Fig. 23 and are all of interest, as they emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and intersect the zero acceleration surface twice. This means they have a decelerated phase, followed by a late-time acceleration as they expand towards $dS_{2+}$.
@@
Figure 22: $X$-$Y$-$Z$ phase space corresponding to Fig. 21, where $w_{x}=-0.6$ and $q=0.847386$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 21 that evolves to the right of the separatrices. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$ and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between between $dS_{2-}$ and $dS_{2+}$, and intersect the zero acceleration surface twice during expansion, meaning they have a decelerated phase and a late-time acceleration.
@@
Figure 23: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 21, where $w_{x}=-0.6$ and $q=0.847386$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.5$, which corresponds to a trajectory in Fig. 21 that evolves to the right of the separatrices, expanding to $dS_{2+}$. Both the flat and open trajectories intersect the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and finally have an accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@ Appendix B $q>3$
The $z$-coordinate of $dS_{1+}$ is proportional to $1/q^{2}$ and the $x$-coordinate is proportional to $1/q$. When $q<3$, $dS_{1+}$ will therefore have larger $z$- and $x$-values than when $q>3$. In particular, when $q>3$, $dS_{1+}$ has an $x$-value of $x<1$. This means the repellor-saddle separatrix will be shorter when $q>3$, as the repelling fixed point is closer to the saddle $dS_{3+}$. We find that when $dS_{1+}$ is a spiral repellor, the separatrix and zero acceleration curve can intersect as the vector field spirals out from $dS_{1+}$, which means the separatrix also spirals out from $dS_{1+}$. However, when $dS_{1+}$ is a repelling or an improper node, the separatrix does not spiral out, and so does not intersect the zero acceleration curve. In the following, we present the two-dimensional and three-dimensional phase spaces where $dS_{1+}$ is a repelling node and an improper node. The color scheme of the fixed points and critical points in the two-dimensional $Z$-$X$ phase spaces is as in Table 4, and the color scheme of the curves is given in Table 5. The color scheme of the fixed points and critical points in the three-dimensional phase spaces is given in Table 8, and the color scheme of the curves and surfaces is as in Table 9.
@@ B.1 Repelling node
Figure 24 shows the $Z$-$X$ phase space when $q>3$ and $dS_{1+}$ is a repelling node. All trajectories are nonsingular, and expand from $dS_{1+}$ at high energy, and tend towards $dS_{2+}$ at late times. The trajectories that do not intersect the red zero acceleration curve always accelerate, and those that cross the red curve twice have a period of deceleration, followed by a late-time acceleration. There is also a limiting case where a trajectory touches the red curve and has a point of zero acceleration, but otherwise accelerates. The trajectories that intersect the zero acceleration curve twice are the models of interest.
@@
Figure 24: $Z$-$X$ phase space referred to in Appendix B.1. Here, the parameters are set to $w_{x}=-0.5$ and $q=5.2$. The stability of $dS_{1+}$ (orange) is a repelling node, and the zero acceleration curve does not intersect the repellor-saddle separatrix. Apart from the separatrix itself which evolves to $dS_{3+}$ (dark blue), all other trajectories expand towards $dS_{2+}$ (magenta) at late times. Trajectories which do not intersect the $a^{\prime\prime}=0$ curve always accelerate, and those that intersect it twice have a decelerated period. One trajectory in the phase space will touch the zero-acceleration curve, but will otherwise accelerate.
@@
Figures 25 and 26 show a first integral surface in the $X$-$Y$-$Z$ phase space corresponding to a trajectory that intersects the zero acceleration curve twice in Fig. 24. Figure 25 shows the trajectories with positive spatial curvature, and Fig. 26 shows expanding flat and open trajectories. All the trajectories in Fig. 25 bounce at least once. Two Einstein points exist at $Y=0$, one of which is a saddle which the CFS passes through, and the other is a center. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, or are cyclic around the center Einstein fixed point. These trajectories are not of interest, as the bouncing trajectories always accelerate, and the cyclic trajectories have a decelerated phase during expansion, but no late-time acceleration. The trajectories that evolve outside the CFS bounce once during a quasi-de Sitter phase, contracting from $dS_{2-}$ then expanding towards $dS_{2+}$. These trajectories intersect the zero acceleration surface twice during expansion, meaning they accelerate though the bounce, becoming flatter as they start to expand, then have a decelerated phase where large scale structure could form, and then have a final late-time accelerated expansion as they approach a late-time cosmological constant represented by $dS_{2+}$. The trajectories outside the CFS are therefore of interest.
@@
Figure 25: $X$-$Y$-$Z$ phase space corresponding to Fig. 24, where $w_{x}=-0.5$ and $q=5.2$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.4$, which corresponds to a trajectory in Fig. 24 that intersects the red zero acceleration curve twice. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and intersect the zero acceleration surface twice during expansion, meaning they have a decelerated phase followed by a late-time acceleration.
@@
Figure 26: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 24, where $w_{x}=-0.5$ and $q=5.2$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.1$ and $Z_{0}=0.4$, which corresponds to a trajectory in Fig. 24 that intersects the red zero acceleration surface twice. Both the flat and open trajectories cross the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.
@@ B.2 Improper node
Figure 27: $Z$-$X$ phase space referred to in Appendix B.2. Here, the parameters are set to $w_{x}=-0.3$ and $q=3.32488$. Here, the stability of $dS_{1+}$ (orange) is an improper node, and the zero acceleration curve does not intersect the repellor-saddle separatrix. Except for the separatrix which expands to $dS_{3+}$ (dark blue), all other trajectories expand towards $dS_{2+}$ (magenta) at late times. Trajectories which do not intersect the $a^{\prime\prime}=0$ curve always accelerate, and those that intersect it twice have a decelerated phase. One trajectory in the phase space will touch the zero-acceleration curve, but will otherwise accelerate.
@@
Figures 28 and 29 show the three-dimensional $X$-$Y$-$Z$ phase spaces corresponding to a trajectory that intersects the zero acceleration curve twice in Fig. 27. Qualitatively, the phase spaces are similar to that of Figs. 25 and 26. Figure 28 shows the trajectories with positive spatial curvature. Two Einstein fixed points exist in the phase space, one of which is a center and the other is a saddle which forms part of the CFS. The trajectories that evolve outside the CFS and bounce once are the models that are qualitatively interesting. These trajectories contract from $dS_{2-}$, then bounce during a quasi de Sitter phase, and finally expand towards $dS_{2+}$. They cross the zero acceleration surface twice during expansion, meaning they accelerate through the bounce, becoming flatter as they start to expand, then have a decelerated period followed by a late-time accelerated expansion. Expanding flat and open trajectories are shown in Fig. 29, and are all of interest. Flat and open trajectories emerge from $dS_{1+}$ and $dS_{4+}$, respectively, and intersect the $a^{\prime\prime}=0$ surface twice. This means they have a decelerated phase where large scale structure could form, and accelerate at late times as they expand towards $dS_{2+}$.
@@
Figure 28: $X$-$Y$-$Z$ phase space corresponding to Fig. 27, where $w_{x}=-0.3$ and $q=3.32488$. The trajectories themselves are plotted on a first integral surface where $X_{0}=0.05$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 27 that intersects the red zero acceleration curve twice. In this figure, only the trajectories with positive spatial curvature are plotted. Two Einstein fixed points (cyan) exist at $Y=0$ on the zero acceleration surface. Trajectories inside the CFS either bounce once between $dS_{2-}$ and $dS_{2+}$, and always accelerate, or are cyclic around an Einstein fixed point and cross the zero acceleration surface once during expansion, meaning they have a decelerated phase but no late-time acceleration. Trajectories outside the CFS bounce once between $dS_{2-}$ and $dS_{2+}$, and intersect the zero acceleration surface twice during expansion, meaning they have a decelerated phase followed by a late-time acceleration.
@@
Figure 29: Expanding flat (green) and open (blue) trajectories in the $X$-$Y$-$Z$ phase space corresponding to Fig. 27, where $w_{x}=-0.3$ and $q=3.32488$. The trajectories are plotted on a first integral surface where $X_{0}=0.05$ and $Z_{0}=0.3$, which corresponds to a trajectory in Fig. 27 that intersects the red zero acceleration curve twice. Both the flat and open trajectories intersect the zero acceleration surface twice, meaning they initially accelerate as they expand from $dS_{1+}$ and $dS_{4+}$, respectively, then have a period of deceleration, and then have a final accelerated phase as they expand toward $dS_{2+}$ at low energy.